A 1.37 kg rock is suspended from a spring that is stretched 0.880 m. What is the spring constant for this spring?
1
Expert's answer
2020-11-30T14:54:53-0500

By the Hooke's Law, we have:



"F=kx,"

here, "F=mg" is the force that acts on the spring, "k" is the spring constant, "x" is the elongation of the sping.

Then, from this formula we can find the spring constant:



"mg=kx,""k=\dfrac{mg}{x}=\dfrac{1.37\ kg\cdot 9.8\ \dfrac{m}{s^2}}{0.880\ m}=15.26\ \dfrac{N}{m}."

Answer:

"k=15.26\ \dfrac{N}{m}."